# DL-Chain: Scalable and Stable Blockchain Sharding with High Concurrency via Dual-Layer Consensus

### v1

@@ 2.2 Blockchain Sharding for Improved Concurrency
Severe Stagnation Issues. We conducted estimations of substantial costs associated with state migrations. Based on the data recorded on the Etherscan, the size of the archive chain data has exceeded 16TB. Even if the node is in pruning mode to synchronize the data, the data volume is over 970GB. We conducted simulations to estimate the time required for a node with a bandwidth of 50Mbps to execute the state migration of Ethereum [2]. The findings reveal that, notwithstanding ledger pruning and utilizing 24 shards, the migration process has extended beyond 100 minutes since 2024, as shown in Figure 2.
@@
Furthermore, estimating the migration procedure, even with the download limited to the state trie for verifying new transactions, consumes more than 25 minutes [5, 3]. Such stagnation issues significantly hamper system functionality, causing a notable decline in performance and user experience.
In DL-Chain, we allow corrupted shards and design a scalable Dual-Layer Consensus to guarantee security without network-wide consensus, achieving high concurrency with more shards. Moreover, we ensure the liveness of shards through rigorous analysis, configurations, and mechanism designs, thus avoiding long-time system stagnation caused by shard migration and providing stable performance.
2.3 Blockchain Systems with Hierarchical Structure
Several approaches have been proposed to enhance scalability by employing a layered, deconstructed design. However, their ideas are different from ours. Prism [9] is a PoW-based blockchain that enhances system scalability by decoupling the processes of transaction packaging and ordering. Blocks satisfying the PoW inequality are randomly categorized into transaction, proposer, and voter blocks. However, Prism is not a sharding system, and the PoW consensus it is based on only provides probabilistic finality and is prone to forking. Moreover, it exclusively ensures consistency in ordering transactions, neglecting to guarantee the correctness of transactions within blocks. Consequently, achieving an accurate ledger mandates a comprehensive traversal of the sorted blocks to eliminate double-spending and duplicate transactions for precise ledger output. In contrast, our system leverages honest nodes in PSs to ensure the validity of transactions and provide deterministic finality.
SSchain [13] introduces a two-layer network structure comprising a root blockchain network as the primary layer and a sharding network as the secondary layer. The root blockchain conducts secondary verification of blocks from shards, mitigating the risk of double-spending. The sharding network contributes to an enhanced overall system throughput. Nevertheless, the system’s scalability encounters limitations due to the need for the root chain maintainer to manage the network-wide ledger. In addition, to ensure security, the system requires a large number of nodes involved in the maintenance of the root chain. In contrast, our system leverages multiple secure finalization committees to ensure security, and nodes do not need to maintain a network-wide ledger.
Pyramid [19] is a novel sharding approach within a two-layer structure to optimize the validation and execution of cross-shard transactions. The system classifies shards into i-shards and b-shards. Each i-shard employs PBFT consensus for intra-shard transactions, while b-shards connect multiple i-shards to process cross-shard transactions independently. However, Pyramid’s requirement for each shard’s honesty and increased number of shards due to the overlapping sharding scheme diminishes system resiliency, only tolerating 16% of malicious nodes in their experiment. In contrast, our system can tolerate 25% of malicious nodes on the network, which is what most shard systems [26, 21, 24] are configured with.
Benzene [11] introduces a PoW-based double-chain sharding scheme where each shard concurrently manages the proposer chain and vote chain. This architecture segregates transaction recording from consensus execution, facilitating cross-shard cooperation verification without impeding independent transaction recording in each shard. The system ensures high fault tolerance, demanding malicious nodes to control over half of the shards to influence vote results. In contrast, our system is based on a BFT-typed consensus protocol that avoids the GPU resources and power consumption associated with the PoW consensus mechanism and can provide deterministic finality.
3 The DL-Chain Model
3.1 Network Model
@@
DL-Chain is deployed on a partial-synchronous Peer to Peer network where there is a known upper bound of delay, denoted as $\delta$, on message transmission delays, which take effect after an unspecified global stabilization time (GST) [35]. As is common with most previous systems [1, 6, 23], messages in DL-Chain are propagated via a gossip protocol.
@@
DL-Chain consists of $N$ nodes, each having a public/private key pair representing its identity when sending messages. Each node belongs to one proposer shard (PS) in the second layer and one finalizer committee (FC) in the first layer simultaneously. There are $C$ FCs, each composed of $K$ disjoint PSs. Hence, the system comprises $C\cdot K$ PSs. Besides, each FC has $n=N/C$ nodes, and each PS has $m=n/K$ nodes.
@@
Each proposer shard, similar to shards in traditional sharding, is responsible for transaction processing. It runs BFT-typed intra-shard consensus to append proposer blocks recording transactions to its proposer chain. Each finalizer committee comprises multiple PSs from the second layer and verifies their proposer blocks. It runs BFT-typed consensus to append finalizer blocks recording hashes of proposer blocks (but not raw transactions) to its finalizer chain.
3.2 Transaction Model
@@
DL-Chain adopts the account/balance model to present the ledger state, the same as existing works [2, 19, 34]. The state (i.e., balance) of a given account is maintained by a single PS in DL-Chain based on the hash of the account address. Each transaction is routed to the corresponding PS based on the related account addresses for processing. Without loss of generality, we will illustrate our system based on normal transfer transactions. However, DL-Chain is also compatible with handling smart contracts, as discussed in Section 6.3.
3.3 Threat Model
@@
In DL-Chain, two categories of nodes exist: honest and malicious. Honest nodes comply with all DL-Chain protocols. For example, they actively respond to consensus messages, refuse to sign blocks containing invalid transactions, and consistently broadcast their signed messages to the entire network. On the other hand, malicious nodes can conduct various types of attacks. In DL-Chain, three main typical attacks can have an additional impact on our system security: (1) silent attack, where they refuse to respond to messages to disrupt consensus; (2) transaction manipulation attack, where they can include invalid transactions into blocks; (3) equivocation attack, where they can send different messages to different nodes. For other typical attacks (e.g., transaction censorship [27], eclipse [18], etc.), there are solutions [14, 30] that are orthogonal to our main designs and can be adopted by DL-Chain (refer to Section 6.3 for more discussion). Besides, it is assumed that malicious nodes cannot forge messages by accessing other nodes’ private keys. We denote the fraction of malicious nodes in the system as $f$, indicating $f\cdot N$ malicious nodes. Similar to existing sharding systems [19, 23, 24], we operate under the assumption of slowly-adaptive adversaries that the distribution of honest and malicious nodes remains static within each epoch (typically one day), and alterations can only occur between epochs.
4 Dual-Layer Consensus
@@
The Dual-Layer Consensus is the backbone component that ensures security while allowing smaller shards for improved concurrency without sacrificing scalability. Specifically, it creates numerous smaller proposer shards for high transaction concurrency and establishes larger finalizer committees to safeguard potentially corrupted PSs. As the number of nodes increases, our system does not rely on network-wide consensus but rather safely establishes additional FCs, each responsible for the security of a subset of PSs, thereby ensuring scalability.
@@
This section describes the basic design of our Dual-Layer Consensus that accomplishes the above goals. However, the system still faces problems on top of this basic architecture. First, the system faces the problem of loss of liveness, which leads to system stagnation and inability to provide stable performance. Second, relying solely on FCs to safeguard the safety of PSs would introduce considerable overhead. Third, since our system may encounter corrupt proposer shards, handling cross-shard transactions poses a unique challenge. Finally, a well-developed system should further guarantee efficiency. We will describe how we deal with these difficult challenges in Section 5.
4.1 Formation of PSs and FCs
@@
DL-Chain runs in fixed periods called epochs with a duration according to the system requirements (e.g., one day for most existing blockchain sharing systems [19, 36, 24]). DL-Chain applies Proof of Stake (PoS) that requires nodes to stake a certain amount of tokens to join the epoch to prevent Sybil attacks, similar to existing works [6, 23]. We assume a trusted beacon chain publically records the identities of these nodes for each epoch, similar to most blockchain sharding [19, 21, 23]. DL-Chain utilizes epoch randomness to assign nodes to PSs randomly to prevent malicious nodes from gathering. The generation of epoch randomness leverages a combination of verifiable random function (VRF) [29] and verifiable delay function (VDF) [10] techniques, as outlined in prior work [6].
@@
Unlike most existing sharding systems, DL-Chain is designed as a Dual-Layered architecture, with each node belonging to a PS and a FC simultaneously. When given a network size $N$ and a fraction of malicious node $f$, the system configures FC size $n$ and PS size $m$, ensuring a negligible probability of failure to obtain the number of PSs securely. Subsequently, the FC identifier of a node is obtained based on the PS identifier of the node (i.e., without loss of generality, a node in $PS_{j}$ belongs to $FC_{\lfloor(j+K-1)/K\rfloor}$ where $K$ represents the number of PSs per FC). In other words, $FC_{i}$ is responsible to safeguard $PS_{j}$ where $j\in[(i-1)K+1,iK]$.
4.2 Strawman Design of Dual-Layer Consensus
@@
We now introduce the strawman design of DL-Chain’s Dual-Layer Consensus. The Dual-Layer Consensus will be more efficient when incorporating the various component designs in Section 5. The basic Dual-Layer Consensus comprises the following three phases:
@@
Block Proposal. In this phase, each PS executes intra-shard consensus to append proposer blocks containing transactions to its proposer chain. As Section 2.1 mentions, we use FBFT [6] as the intra-shard consensus. However, we derive a different quorum size $m/2+1$ for a PS of size $m$ (the rationale is shown in Section 5.2. PSs cannot guarantee safety on their own, that is why they need FCs’ assistance). Like most systems, honest nodes verify raw transactions before voting for proposer blocks. Nevertheless, a PS may be corrupted due to its smaller size. Hence, each honest node must broadcast the proposer block (after the designs in Section 5.2, only header is required) it voted for to the corresponding FC for verification and finalization later. This design guarantees the block will be broadcast to the FC if it passes the intra-shard consensus and is voted by at least one honest node.
@@
Next, nodes within the FC verify the validity of transactions in proposer blocks. Our basic design presumes FC nodes keep track of all corresponding PS states to verify each raw transaction. A more efficient design refining this process is detailed in Section 5.2. After validation, each FC node retains valid proposer blocks in its memory.
@@
Block Finalization. In this phase, each FC performs standard BFT-typed consensus (e.g., FBFT, as mentioned in Section 2.1, tolerating $<n/3$ malicious nodes with $12n/3+$ 1 quorum size) to produce finalizer blocks containing hashes of valid proposer blocks. Honest FC nodes verify the validity of proposer blocks recorded in finalizer blocks. They can use cached verification results from the previous phase to speed up. Additionally, nodes must ensure that finalized proposer blocks do not form forks for consistent finalization. After passing the FC’s consensus, honest nodes broadcast a finalizer block to the corresponding PSs. This allows PSs to confirm the execution of finalized proposer blocks and update the ledger state.
@@
Balance of Scalability and Security. The Dual-Layer Consensus relies on FCs for security. We ensure that FC sizes are minimal, preserving a high likelihood of being secure to balance security and scalability. Section 6.1 details the complete derivation and proof.
5 Enhanced Designs for DL-Chain
@@
Based on the Dual-Layer Consensus, in this section, we introduce the core and unique designs of DL-Chain to achieve stable performance, low overhead, secure cross-shard transaction handling, and high efficiency. First, to address the system stagnation issue and provide stable performance, we design a cross-layer view change mechanism and ensure the quorum size within PSs (even if corrupted) can be reached so that each PS will not lose liveness. This is introduced in Section 5.1. Second, to reduce cross-layer communication overhead while ensuring safety, we rely on consensus within PSs to ensure the validity of transactions and then utilize FCs to resolve the forking issues of PSs. Therefore, only headers are transmitted across two layers for low overhead. This part will be explained in Section 5.2. Third, as will be described in Section 5.3, to guarantee cross-shard transaction security, we require the finalization of cross-shard transactions to depend on secure FCs. Finally, to achieve higher system efficiency, we design a pipeline mechanism in which PSs and FCs produce blocks concurrently, which will be discussed in Section 5.4.
5.1 Stagnation Prevention
@@
In the first scenario, a PS’s malicious leader stops block production. In most BFT-typed consensus protocols [12, 35, 6], leaders package transactions into blocks to be voted for. Without block production, the consensus process cannot proceed. Although the consensus typically has a view change mechanism to substitute a malicious leader, a corrupted PS has too many malicious nodes (exceeding the BFT-typed consensus’s tolerance threshold, $1/3$) to replace a malicious leader independently. Hence, we propose the cross-layer view change mechanism using FCs to replace the malicious PS leader. The mechanism includes these steps:
Complaint for the Leader. Suppose a node has not accepted blocks from the leader within a timeout (initialized when the last consensus is reached, maintained by the exponential back-off mechanism, the same as most existing works [12, 35, 19]). In that case, it suspects leaders of remaining silent and initiates the cross-layer view change by sending Complain message to its PS and its FC. This message should contain the suspected leader’s identifier and the reasons, such as a lack of block proposal.
@@
Discussion of Cross-Layer View Change. The security of this mechanism depends on FCs, which is the same as our Dual-Layer Consensus. Fortunately, we guarantee the security of FCs with high probability after rigorous theoretical derivations and proof, as shown in Section 6.1. Besides, this view change mechanism can also replace malicious leaders that launch other attacks (e.g., equivocation and transaction manipulation attacks). Specifically, honest nodes must provide two blocks of the same height signed by the same leader to prove an equivocation attack or a block containing an incorrect transaction to prove a transaction manipulation attack (after the designs in next Section 5.2, only block headers are required to detect equivocation, and the transaction validity can be guaranteed inside each PS). FCs then can verify the feasibility of Complain messages.
@@
In the second scenario, the PS has fewer honest nodes than the quorum size. In this case, malicious nodes can collectively remain silent to prevent valid blocks from collecting quorum size of votes, thus obstructing consensus. To address this issue, we aim to ensure that honest nodes are more than or equal to the quorum size with a high probability. We accomplish this through rigorous theoretical derivations and configuration, as detailed in Section 6.1. As a result, even if malicious nodes are silent, they can’t prevent a valid proposer block from getting enough votes from honest nodes and passing consensus.
5.2 Overhead Reduction

### v2

@@ II-B Blockchain Sharding for Improved Concurrency
We conducted estimations of substantial costs associated with state migrations. Based on the data recorded on the Etherscan, the size of the archive chain data has exceeded 16TB. Even if the node is in pruning mode to synchronize the data, the data volume is over 970GB. We conducted simulations to estimate the time required for a node with a bandwidth of 50Mbps to execute the state migration of Ethereum [eth]. The findings reveal that, notwithstanding ledger pruning and utilizing 24 shards, the migration process has extended beyond 100 minutes since 2024, as shown in Figure 2.
@@
Furthermore, estimating the migration procedure, even with the download limited to the state trie for verifying new transactions, consumes more than 25 minutes [stateTrie, headState]. Such stagnation issues significantly hamper system functionality, causing a notable decline in performance and user experience.
Our Positioning. In StableShard, we also permit corrupted proposer shards, but we differ from the above lines in a key way: we avoid recovery-after-failure altogether. Instead of respawning/resizing shards after liveness breaks (GearBox), or replacing/taking over corrupted shards after detection (CoChain), StableShard keeps shard membership and states intact and maintains progress in-place through PS-FC cooperation.
Concretely, StableShard co-designs (i) a complementary division of labor and (ii) matched parameters: PSs guarantee transaction validity locally via quorum voting, enabling FCs to remain stateless and finalize using only headers for fork-resolution and finality; FCs additionally maintain stagnation-free liveness for faulty PSs via a cross-layer view-change protocol, so a corrupted PS does not trigger reshuffling or state migration. By jointly fine-tuning committee sizes and quorum thresholds, StableShard supports many small PSs (tolerating Byzantine fraction approaching <1/2<1/2) while allowing multiple FCs to coexist securely (each <1/3<1/3), thus achieving high concurrency with stable and scalable performance.
III The StableShard Model
III-A Network Model
@@
StableShard is deployed on a partial-synchronous Peer to Peer network, in which there is a known bound $\delta$ and an unknown Global Stabilization Time (GST). After GST, the network becomes synchronous, and all transmissions between two honest nodes arrive within time $\delta$ [hotstuff]. As is common with most previous systems [bitcoin, Harmony, li2022jenga], messages in StableShard are propagated via a gossip protocol.
@@
StableShard consists of $N$ nodes, each having a public/private key pair representing its identity when sending messages. Each node belongs to one proposer shard (PS) in the second layer and one finalizer committee (FC) in the first layer simultaneously. There are $C$ FCs, each composed of $K$ disjoint PSs. Hence, the system comprises $C\cdot K$ PSs. Besides, each FC has $n=N/C$ nodes, and each PS has $m=n/K$ nodes.
@@
Each proposer shard, similar to shards in traditional sharding, is responsible for transaction processing. It runs BFT-typed intra-shard consensus to append proposer blocks recording transactions to its proposer chain. Each finalizer committee comprises multiple PSs from the second layer and verifies their proposer blocks. It runs BFT-typed consensus to append finalizer blocks recording hashes of proposer blocks (but not raw transactions) to its finalizer chain. The consensus running in each FC is lightweight, as it does not have to handle heavy transactions.
III-B Transaction Model
@@
StableShard adopts the account/balance model to present the ledger state, the same as existing works [eth, pyramid, monoxide]. The state (i.e., balance) of a given account is maintained by a single PS in StableShard based on the hash of the account address. Each transaction is routed to the corresponding PS based on the related account addresses for processing. Without loss of generality, we will illustrate our system based on normal transfer transactions. However, StableShard is also compatible with handling smart contracts, as discussed in Section VI-C.
III-C Threat Model
@@
In StableShard, two categories of nodes exist: honest and malicious. Honest nodes comply with all StableShard protocols. For example, they actively respond to consensus messages, refuse to sign blocks containing invalid transactions, and consistently broadcast their signed messages to the entire network. On the other hand, malicious nodes can conduct various types of attacks. In StableShard, three main typical attacks can have an additional impact on our system security: (1) silent attack, where they refuse to respond to messages to disrupt consensus; (2) transaction manipulation attack, where they can include invalid transactions into blocks; (3) equivocation attack, where they can send different messages to different nodes. For other typical attacks (e.g., transaction censorship [censorship], eclipse [eclipse], etc.), there are solutions [redBelly, honeybadger] that are orthogonal to our main designs and can be adopted by StableShard (refer to Section VI-C for more discussion). Besides, it is assumed that malicious nodes cannot forge messages by accessing other nodes’ private keys. We denote the fraction of malicious nodes in the system as $f$, indicating $f\cdot N$ malicious nodes. Similar to existing sharding systems [pyramid, li2022jenga, cochain], we operate under the assumption of slowly-adaptive adversaries that the distribution of honest and malicious nodes remains static within each epoch (typically one day), and alterations can only occur between epochs.
IV Division-of-Labor Collaboration Protocol
@@
The PS-FC Division-of-Labor Collaboration (DoLC) protocol is the backbone component that ensures security while allowing smaller shards for improved concurrency without sacrificing scalability. Specifically, it creates numerous smaller proposer shards for high transaction concurrency and establishes larger finalizer committees to safeguard potentially corrupted PSs. As the number of nodes increases, our system does not rely on network-wide consensus but rather safely establishes additional FCs, each responsible for the security of a subset of PSs, thereby ensuring scalability.
@@
This section describes the basic design of our DoLC protocol that accomplishes the above goals. However, the system still faces problems on top of this basic architecture. First, the system faces the problem of loss of liveness, which leads to system stagnation and inability to provide stable performance. Second, relying solely on FCs to safeguard the safety of PSs would introduce considerable overhead. Third, since our system may encounter corrupt proposer shards, handling cross-shard transactions poses a unique challenge. Finally, a well-developed system should further guarantee efficiency. We will describe how we deal with these difficult challenges in Section V.
IV-A Formation of PSs and FCs
@@
StableShard runs in fixed periods called epochs with a duration according to the system requirements (e.g., one day for most existing blockchain sharing systems [pyramid, rapidchain, cochain]). StableShard applies Proof of Stake (PoS) that requires nodes to stake a certain amount of tokens to join the epoch to prevent Sybil attacks, similar to existing works [Harmony, li2022jenga]. We assume a trusted beacon chain publically records the identities of these nodes for each epoch, similar to most blockchain sharding [pyramid, omniledger, li2022jenga]. StableShard utilizes epoch randomness to assign nodes to PSs randomly to prevent malicious nodes from gathering. The generation of epoch randomness leverages a combination of verifiable random function (VRF) [micali1999verifiable] and verifiable delay function (VDF) [boneh2018verifiable] techniques, as outlined in prior work [Harmony].
@@
Unlike most existing sharding systems, StableShard is designed as a division-of-labor and collaboration between PSs and FCs, with each node belonging to a PS and a FC simultaneously. When given a network size $N$ and a fraction of malicious node $f$, the system configures FC size $n$ and PS size $m$, ensuring a negligible probability of failure to obtain the number of PSs securely. Subsequently, the FC identifier of a node is obtained based on the PS identifier of the node (i.e., without loss of generality, a node in $PS_{j}$ belongs to $FC_{\lfloor(j+K-1)/K\rfloor}$ where $K$ represents the number of PSs per FC). In other words, $FC_{i}$ is responsible to safeguard $PS_{j}$ where $j\in[(i-1)K+1,iK]$.
IV-B Strawman Design of DoLC protocol
@@
We now introduce the strawman design of StableShard’s DoLC protocol. The DoLC protocol will be more efficient when incorporating the various component designs in Section V. The basic DoLC protocol comprises the following three phases:
@@
Block Proposal. In this phase, each PS executes intra-shard consensus to append proposer blocks containing transactions to its proposer chain. As Section II-A mentions, we use FBFT [Harmony] as the intra-shard consensus. However, we derive a different quorum size $m/2+1$ for a PS of size $m$ (the rationale is shown in Section V-B. PSs cannot guarantee safety on their own, that is why they need FCs’ assistance). Like most systems, honest nodes verify raw transactions before voting for proposer blocks. Nevertheless, a PS may be corrupted due to its smaller size. Hence, each honest node must broadcast the proposer block (after the designs in Section V-B, only header is required) it voted for to the corresponding FC for verification and finalization later. This design guarantees the block will be broadcast to the FC if it passes the intra-shard consensus and is voted by at least one honest node.
@@
Next, nodes within the FC verify the validity of transactions in proposer blocks. Our basic design presumes FC nodes keep track of all corresponding PS states to verify each raw transaction. A more efficient design refining this process is detailed in Section V-B. After validation, each FC node retains valid proposer blocks in its memory.
@@
Block Finalization and Fork Selection. In this phase, each FC performs standard BFT-typed consensus (e.g., FBFT, as mentioned in Section II-A, tolerating <n/3<n/3 malicious nodes with $2n/3+1$ quorum size) to produce finalizer blocks containing hashes of valid proposer blocks. Honest FC nodes verify the validity of proposer blocks recorded in finalizer blocks. They can use cached verification results from the previous phase to speed up. Additionally, each FC assists its corresponding PSs in fork selection through intra-shard consensus, as will be explained later. After passing the FC’s consensus, honest nodes broadcast a finalizer block to the corresponding PSs. This allows PSs to confirm the execution of finalized proposer blocks and update the ledger state.
The process by which FCs assist PSs in fork selection is critical, as it determines which fork from a potentially corrupted PS will ultimately be finalized by the FC. In partial-synchronous networks, deterministically detecting forks is impossible. Fortunately, StableShard does not require deterministic fork detection. Within any FC, the leader follows two primary rules to select a fork from its corresponding PSs. First, the selected fork must be a continuation of the last fork finalized by the FC to prevent conflicts. Second, there may be multiple forks that all extend from the same fork previously finalized by the FC. Among these, the leader selects the longest fork from its own perspective and proceeds to finalize the latest block on this fork, along with other transactions, through intra-shard consensus. Honest nodes within the FC, when participating in consensus, will verify whether the fork selected by the leader is indeed a continuation of the last fork finalized by the FC and whether the leader has selected only one fork (to prevent equivocation attacks). However, these honest nodes do not judge whether the leader has selected the longest fork, as different nodes may have inconsistent views of the network state in a partial-synchronous environment. This fork selection mechanism relies on the intra-shard consensus of the FC. Therefore, when the intra-shard consensus within the FC is secure, this mechanism ensures the security of fork selection.
@@
Balance of Scalability and Security. The DoLC protocol relies on FCs for security. We ensure that FC sizes are minimal, preserving an extremely high likelihood of being secure to balance security and scalability. Section VI-A details the complete derivation and proof.
V Enhanced Designs for StableShard
@@
Based on the basic DoLC protocol, in this section, we introduce the core and unique designs of StableShard to achieve stable performance, low overhead, secure cross-shard transaction handling, and high efficiency. First, to address the system stagnation issue and provide stable performance, we design a cross-layer view change mechanism and ensure the quorum size within PSs (even if corrupted) can be reached so that each PS will not lose liveness. This is introduced in Section V-A. Second, to reduce cross-layer communication overhead while ensuring safety, we rely on consensus within PSs to ensure the validity of transactions and then utilize FCs to resolve the forking issues of PSs. Therefore, only headers are transmitted across two layers for low overhead. This part will be explained in Section V-B. Third, as will be described in Section V-C, to guarantee cross-shard transaction security, we require the finalization of cross-shard transactions to depend on secure FCs. Finally, to achieve higher system efficiency, we design a pipeline mechanism in which PSs and FCs produce blocks concurrently, which will be discussed in Section V-D.
V-A Stagnation Prevention
@@
In the first scenario, a PS’s malicious leader stops block production. In most BFT-typed consensus protocols [pbft, hotstuff, Harmony], leaders package transactions into blocks to be voted for. Without block production, the consensus process cannot proceed. Although the consensus typically has a view change mechanism to substitute a malicious leader, a corrupted PS has too many malicious nodes (exceeding the BFT-typed consensus’s tolerance threshold, $1/3$) to replace a malicious leader independently. Hence, we propose the cross-layer view change mechanism using FCs to replace the malicious PS leader. The mechanism includes these steps:
Complaint for the Leader. Suppose a node has not accepted blocks from the leader within a timeout (mentioned in Section II-A). In that case, it suspects leaders of remaining silent and initiates the cross-layer view change by sending Complain message to its PS and its FC. This message should contain the suspected leader’s identifier and the reasons, such as a lack of block proposal.
@@
Discussion of Cross-Layer View Change. The security of this mechanism depends on the FCs, which is the same as our DoLC protocol. Fortunately, we guarantee the security of FCs (each with $1<3$ fraction of malicious nodes) with high probability after rigorous theoretical derivations and proof, as shown in Section VI-A. Besides, this view change mechanism can also replace malicious leaders that launch other attacks (e.g., equivocation and transaction manipulation attacks). Specifically, honest nodes must provide two blocks of the same height signed by the same leader to prove an equivocation attack or a block containing an incorrect transaction to prove a transaction manipulation attack (after the designs in next Section V-B, only block headers are required to detect equivocation, and the transaction validity can be guaranteed inside each PS). FCs then can verify the feasibility of Complain messages.
The above mechanism is secure both before and after GST, provided that the intra-shard consensus within the FC is secure. After GST, its security is unequivocal because the network transitions to a synchronous state, and timeouts stabilize. Before GST, the mechanism remains secure because it relies on the assistance of the secure FC to perform cross-layer view changes for its corresponding PSs. The intra-shard consensus within each FC is no different from existing BFT-type consensus designed for partial-synchronous networks, which are equipped with dynamically increasing timeout mechanisms and robust view change protocols (as detailed in Section II-A). Therefore, even before GST, when the FC is secure (i.e., the proportion of malicious nodes <1/3), its intra-shard consensus can guarantee both safety and liveness (liveness may be temporarily affected due to asynchrony). As a result, each FC can securely assist PSs in cross-layer view change, even before GST.
@@
In the second scenario, the PS has fewer honest nodes than the quorum size. In this case, malicious nodes can collectively remain silent to prevent valid blocks from collecting quorum size of votes, thus obstructing consensus. To address this issue, we aim to ensure that honest nodes are more than or equal to the quorum size with a high probability. We accomplish this through rigorous theoretical derivations and parameterization, as detailed in Section VI-A. As a result, even if malicious nodes are silent, they cannot prevent a valid proposer block from getting enough votes from honest nodes and passing consensus.
V-B Overhead Reduction